OpenStudy (anonymous):

x - 2 > 2x + 1 or -2x - 2 < -10

OpenStudy (anonymous):

x - 2 \[\le\]2x + 1 or -2x - 2 < -10

OpenStudy (aroub):

The second one is the correct question?

OpenStudy (anonymous):

its all one line

OpenStudy (aroub):

x-2<=2x+1 or -2x-2<-10 -2x+x<=2+1 -2x<-8 -x<=3 x>4 x>=-3 => x>=-3 or x>4

OpenStudy (anonymous):

will you please draw a graph?

OpenStudy (aroub):

Sure!

OpenStudy (aroub):



OpenStudy (aroub):

Since its a disjunction the graph will become:

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

how wud u graph x.2 and x-7?

OpenStudy (anonymous):

x>2

OpenStudy (anonymous):

x\[\le\]-7

OpenStudy (aroub):